# Towards an Accessible, Noninvasive Micronutrient Status Assessment Method: A Comprehensive Review of Existing Techniques

### v1

@@ 3.6 Biofluid Analysis Limitations
Prediction from clinical health data is able to combine and analyze a large breadth of features, but this data is often insufficient for micronutrition. The lack of micronutrition data availability poses a grand limitation for the ability to make strides in analytic techniques with AI/ML. As a result, Brown et al. [21] and others urge for more micronutrition data. Additionally, the insights provided by these solutions are limited to a small set of micronutrients and/or indicate only a binary deficiency status, rather than a continuous one.
4 Physiological Analysis Methods for Micronutrient Status Assessment
This section outlines methods that assess micronutrient status by detecting associated physiological features. Physiological assessment is just as important as biofluid assessment for micronutrition, with many arguing that the two should be considered jointly. Despite this, physiological analysis receives significantly less attention in emerging research.
4.1 Clinical Evaluation and Physical Examination
An important component when clinically diagnosing a nutritional imbalance is analyzing the medical history of the patient and conducting a physical exam [142, 68]. When considering a patient’s medical history, the assessor looks for factors that influence nutritional intake, needs, and absorption like medication, diseases, impairments in physical functions, and operations. A nutrition-focused physical exam (NFPE) is "an objective method of detecting clinical signs and symptoms of nutritional deficiencies" [142]. A typical NFPE examines muscle mass, fat stores, fluid retention, and specific nutrient deficiencies. The first three involve assessment of an individual’s “bulk and tone”, “bone visibility”, and a “spongy” feel. Muscle strength and function is also assessed by handgrip strength testing. A standalone study on the NFPE goes into more detail on its methodology and efficacy [68]. Common regions of interest (ROIs) for nutrient deficiencies are the skin, nails, scalp, eyes, and mouth. Different symptoms in these regions have been shown to have associations with nutrient deficiencies (mostly in micronutrients). For example pallor, or white coloring on the nails, indicates a possible iron, protein, and vitamin B12 deficiency. Such symptoms and their relations to individual deficiencies are summarized in Tables LABEL:tab:phys-vitamins-water-soluble, LABEL:tab:phys-vitamins-fat-soluble, and LABEL:tab:phys-minerals. Additionally, some regions such as the skin, mouth, hair, and tongue have high cellular turnover, meaning symptoms of deficiency can manifest relatively quickly in these regions. This characteristic could be key to the early detection of clinical deficiencies. Due to its subjectivity, an expert is often necessary to perform a NFPE. It is also important to repeat the exam periodically to take changes over time into account.
4.1.1 Dietary History
Nutritional status is commonly determined by the balance between food intake and energy expenditure, as reported by dietary history [142]. Dietary history is another component of assessment, and possibly one of the most flawed. It looks at a person’s intake patterns, habits, and other variables such as culture, allergies, diet, and more. To analyze this balance, the assessor must have an idea of an individual’s nutritional requirements, as well as their dietary intake. The gold standard for determining the former is indirect calorimetry, which provides insights into a person’s energy expenditure, and therefore primarily macronutrient use. The latter half of the balance, dietary intake, attempts to quantify the amount of nutrients in one’s diet. The authors of one study suggest that “the assessment of macronutrients (fat, carbohydrates, and proteins) is as important as the assessment of micronutrients (vitamins, trace elements)” [142]. While this should be the case, we see that micronutrients are often an afterthought in dietary logging. Because EDRs are not one-size-fits-all, estimates of nutrient requirements are useful to paint a fuller picture of an individual diet. Although indirect calorimetry and formulas considering energy expenditure have been shown to be effective for macronutrient requirements estimation, no such method exists for micronutrients [142]. Macronutrient intake can be correlated with micronutrient intake, however this is not fully indicative of micronutrient intake.
AI and ML have become prominent tools in gathering and analyzing dietary history data. Macronutrient status monitoring in particular has used AI to estimate energy expenditure from physical activity data [146]. Various work has used AI and dietary pattern data to predict the risk of diseases, such as breast cancer, colorectal cancer, and cardiometabolic diseases. This technology can also help study the effects of nutrient supplementation on disease treatment and prevention, or examine relationships between mineral levels and cardiovascular disease, diabetes, and schizophrenia. Dietary recommendations can be made on an individual level using information from nutritional databases and clinical expertise [146, 36]. Various types of data can be useful to a nutrition-focused AI model, even social media. As reported by Côté and Lamarche [36], several studies have gathered and analyzed nutritional behavior data from Twitter and Reddit to uncover dietary correlations and understand geographical and demographic differences in diet. Computer vision has also been used to identify and estimate nutrient intake from consumed food and extract eating behaviors from video [36, 146, 72]. Some researchers have designed smart neckbands that utilize AI and ML to track dietary intake and provide nutritional insights [129, 30, 83].
Figure 4: Having an interdependent perspective on all aspects of nutrition can lead to a better understanding of its health effects. Used with permission from [139].
4.2 Remote Nutrition Assessment
Since the COVID-19 pandemic, there has been a growing push for assessment methods that don’t require a clinician to be physically-present [68, 109, 155]. These remote assessments use telehealth technology to carry out clinical evaluations and NFPEs. Many components of the NFPE remain the same in the remote setting, with increased reliance on activity tracking devices when assessing a patient’s functional status. Observing a patient’s living space during a telehealth call is another useful way a way to gain insights into nutritional intake, food availability, and quality of life. The visual inspection of ROIs and the analysis of health data is critical to conducting a remote NFPE. A video call or a collection of photos is especially helpful, as the nutritionist can visually inspect ROIs like they would in-person. It is recommended to start with the oral cavity since it has many early signs of deficiency [68]. Health tracking devices can be useful for a remote exam, providing the practitioner with biometric and physical activity data that can guide a telehealth visit and/or nutritional intervention. Consumer-level scales, bioelectrical impedance analysis (BIA) devices, blood glucose monitors, blood pressure monitors, sleep trackers, and activity trackers are readily available as sources of these data.
Verbal questioning can also be a valuable source of information for remote nutrition assessments, especially for less tech-savvy patients or when photo sharing or video calls are not possible. Researchers recommend asking the patient a series of specific questions to guide them through a virtual head-to-toe examination [109, 155]. Chizmar and Lewis [33] assessed the correlation between patient responses to verbal questions from a clinician and “physical signs of malnutrition” to validate assessments of nutrition without a physical exam. A set of four questions about eye and rib appearance, percieved skinniness, and intake changes from swelling were significantly associated with a diagnosis of malnutrition. These and other questions were further associated with other individual factors of malnutrition. As to be expected, this study has few specific implications for micronutrition, but it’s findings are nonetheless useful for the diagnosis of nutrition as a whole. It should be noted that of these 30 patients, 22 were white and 28 were male. This is critical, as female and non-white patients are disproportionately affected by the burden of malnutrition [134, 80].
In remote dietary assessment are app-based, most new innovations are app-based [109]. These methods can act as intake loggers: users can input their dietary history, and nutritional metrics can be calculated from food and nutrient databases. There is also potential for remote dietary assessment apps to connect with health-tracking devices that can provide additional context to a patient’s dietary history.
4.3 Physiological Sensing
Physiological sensing may be useful for the automatic detection of physiological symptoms of deficiency (Section 2.3), aiding NFPE assessments. King et al. [87] reviews techniques for yielding health insights from multiple sensors and their applications. Applications here focus on aspects of health that are not directly related to nutrition, however the authors suggest several criteria for practical wearable monitoring, such as being "noninvasive, intuitive to use, reliable, and provide relevant feedback to the wearer" [87]. Witt et al. [184] describes interesting ways that raw sensor data from wearables can be used to explore health. Different sensors such as photoplethysmogram (PPG), ECG, accelerometer, EDA, and those collecting temperature and physical activity can give valuable insights into human physiology that can be associated with micronutrient imbalance. Some examples are insights into HRV, sleep quality, stress levels, and behavioral patterns. While not explicitly wearables-focused, Yokus and Daniele [187] provide some useful considerations for wearable devices that should be applied to future micronutrient assessing devices.
4.3.1 Electrodermal Activity (EDA)
EDA measures the change in skin conductance caused by sweat, which is an indicator of nervous system arousal [136]. EDA is a common sensing modality in wearables, but we feel it is unexplored in the context of micronutrition. We review the potential of EDA as a measure of the nervous system, which can be associated with micronutrient status (Section 2.3).
All eccrine sweat glands activate upon sympathetic nervous system arousal [136]. As a signal, EDA is based on how “sweat gland activity modulates the conductance of an applied current” [136]. EDA signal analysis can yield several interesting metrics, such as skin conductance responses, nonspecific skin conductance responses, and skin conductance levels. These give insight into sympathetic response and conductance measured. One study shows the potential of EDA, when used together with HRV, to yield insights into nervous system functionality [57]. The approach combines HRV point process models and EDA spectral analysis to better detect autonomic nervous system response to cold-pressor and emotional response tests. They report an accuracy of 73.08% for “automatic emotional valence recognition”.
EDA’s spectral component varies based on activity level [136]. Efforts have been made to separate high and low frequencies of EDA to isolate sensitive responses. Fingerprinting of EDA signals is an interesting concept, with one paper proposing a spectrogram-inspired multidimensional representation of EDA signals [28]. This approach allows the representation to emphasize finer signal changes, helping to distinguish between arousal levels and stress-eliciting environments.
4.3.2 Photoplethysmography (PPG)
Similarly to EDA, PPG has a lack of attention in its potential to derive micronutritional insights, yet it is even more prevalent in wearable and accessible devices. It is commonly found in pulse-oximiters, smartwatches, and other health sensors. PPG is a continuous signal of cardiac function obtained by measuring the light transmitted through or reflected by the skin (Fig. 5). As blood pulses, it is possible to observe subtle color changes corresponding to the pulse rate (PR), analogous to heart rate (HR). From PR, we can derive pulse rate variability (PRV), which can provide valuable insights into the autonomic nervous system (ANS) activity [99]. PPG can further derive HRV and blood pressure, thus giving insight into bodily functions. Other works have used raw PPG data to estimate BP via ML [114], and DL [127]. Some notable works that apply PPG include an open-source method for remote PPG using a consumer-grade webcam [175]. Smartphone cameras have also been used to conduct PPG through the fingertip, with the smartphone’s flash as a light source [59].
4.3.3 Optical Sensors
@@
McDuff [110] reviews how camera sensors can be used for non-invasive physiological measurement, which can allow for nutritional insights. The analysis of motion artifacts can reveal minute subtleties in body motion over time that are caused by various physiological mechanisms (e.g. breathing). Also, camera sensors can measure the intensity and wavelength of light absorbed and reflected by our bodies, especially skin (Fig. 5). Differences in measured light over time can be observed and associated with physiological signals (e.g. heartbeat) and status (e.g. low blood oxygen saturation).
@@
Figure 5: As light is applied to the skin, it is absorbed and reflected in different ways depending on its wavelength. Used with permission from [110].
@@
The type of optical sensor has a large impact on the signals that can be derived from it [110]. RGB cameras are found in most smartphones and, therefore, are the most prevalent. These operate largely in the visible spectrum of light (400 to 700 nm wavelengths), but they can often detect some light in the NIR range. NIR cameras are able to detect light in the 700 to 1000 nm range, and thermal cameras can go fully into the infrared spectrum of 2000 to 14000 nm. Thermal cameras, as their name suggest, can provide unique information over other sensors such as body temperature and sweat gland activation. However, this comes with a higher cost and lower resolution. Finally, multi and hyper-spectral cameras allow for the measurement of multiple wavelengths of light at once. This can also be achieved by combining signals from multiple sensors.
@@
Cameras have been extensively applied to the measurement of physiological vital signs [110, 111, 100, 117, 175, 59]. One survey describes 5 physiological vital signs where optical sensor measurements been applied: pulmonary activity, EDA, blood oxygen saturation, glucose status, and cardiac activity [110]. PPG signal caputure has been achieved with RGB cameras. NIR cameras are also useful for PPG in use cases like sleep monitoring that require environments with low visible light [110]. When measuring blood oxygen saturation, multi or hyper-spectral cameras are preferred. Measuring multiple wavelengths of light at once benefits the simultaneous measurement of oxy and deoxy-hemoglobin, and therefore blood oxygen saturation. Optical sensor-based techniques could be use to detect and investigate other physiological symptoms in the fingertip [62, 179, 173], mouth [183, 186], and eyes [164, 128, 93, 79].
@@
Optical sensors have been successfully applied to noninvasively assess a major protein, hemoglobin. Hemoglobin allows for oxygen transport in red blood cells and is produced by iron, vitamin B9, and vitamin B12 [161]. A deficiency in these micronutrients can result in lowered hemoglobin, which manifests as anemia. A review of the state of the art emphasizes the need for an affordable and accessible method of hemoglobin measurement, which can be realized with commodity smartphone cameras [63]. Various work has used smartphone cameras as a way to estimate hemoglobin status noninvasively [63, 179, 62, 164]. They report that PPG signals derived from the fingertip and conjunctiva (skin behind the lower eyelids) under NIR light in 1070 and 850 nm wavelengths contain the most critical features to hemoglobin estimation. Wang et al. [179] had users place their fingertip directly onto a smartphone camera while it recorded video to determine hemoglobin status. Analysis found that the blue spectra of plasma was the most important for protein composition, and therefore hemoglobin estimation. Classifiers were built to identify hemoglobin status as “normal" or “anemic" in the context of demographic averages. Gold-standard blood hemoglobin measurements (Masimo Pronto optical device) were used as ground-truth. Because camera-based assessment can exhibit bias with differences in skin color, the authors carefully considered and reported the demographics of their study participants in addition to controlling for light absorption by skin tissue during experiments. A later paper devised a similar system, applying artificial neural networks (ANNs) and using only the smartphone’s flash as a light source [62]. By combining frames within the videos, the researchers were able to identify regions with high variation in predicted levels.
@@
The use of the conjunctiva as an ROI is gaining popularity due to the fact that it, like the fingernail bed and palmar creases, has no melanin and is devoid of “epidermis, dermis or subcutaneous fat which could impede the transmission of light” to deeper vascular layers [164]. This means the blood vessels are easier to analyze and there may be less bias due to skin color. These properties have been leveraged by Suner et al. [164] to estimate hemoglobin concentration and screen for anemia using smartphone images of the conjunctiva. Spectral super-resolution (SSR) has been introduced to measure blood hemoglobin levels [128]. This method is based on a wealth of existing research reconstructing hyperspectral images from RGB signals. Statistical learning is applied to approximate a hyperspectral image of the conjunctiva from a simple smartphone camera image. Using the hyperspectral data, the hemoglobin content in blood can be computed more effectively than the RGB data alone. However, no analysis of skin color differences was conducted.
4.4 Physiological Analysis Limitations
The general limitation of physiology-based micronutrient assessments is that physiological symptoms are often only shown to be associated with deficiency, and not proven to be causally-related. This limitation could stem from the intricacy of micronutrients and the cost and complexity of running a clinical micronutrition study, which results in a lack of micronutritional data to analyze and mine for new insights. This echoes the limitation of a lack of micronutritional-related data that we found in biofluids analysis. The largest form of physiological micronutritional data are dietary surveys, which often prove to be unreliable [142]. This review finds a lack of technological advancements in physiology-based micronutrient assessments as compared to biofluid-based ones.
Physiological assessment of micronutrient status from the clinical perspective can be summarized by nutrition-focused physical examinations and dietary logs. Physical exams are subjective, qualitative, and require a trained professional. Dietary logs are the most popular and accessible. However, they provide sparse micronutrient information, rely heavily on self-reporting, and do not provide an objective measure for dietary intake in free-living individuals [142]. These accuracy and bias issues are especially a concern when estimating micronutrient intake, where precision is key. AI and ML tools for dietary logging do not adequately address it’s fundamental shortcomings, and face additional problems of low explainability, a high potential for induced bias, a general lack of performance comparisons, and a persistent debate over where these methods should exist in the patient-clinician relationship [36, 146].
The remote approach to nutrition assessment relies on patient input and cooperation, as well as the patient’s own eye for detail, more so than non-remote approaches [109, 155]. Remote nutrition assessments also inherit common telehealth problems, including wireless signal and camera/microphone quality issues. The accuracy of health tracking devices used to aid these assessments varies and they must be referenced consistently over time to provide meaningful information to the clinician. Remote dietary assessment methods are not immune from the limitations in dietary intake logging described previously. Even in light of these limitations, remote assessment methods can be as effective as traditional methods, while being more accessible [109].
Physiological sensing is the most accessible and noninvasive avenue for micronutrient status assessment. However, for physiological sensing to be accurate, the detected signals and symptoms must be placed in a clinical context. Otherwise, the thin line between association and causation is at risk of being blurred. We also note that current physiological analysis approaches fail to fully explore the variety of physiological symptoms of micronutrient imbalance (Section 2.3).
@@ 4.4 Physiological Analysis Limitations
Techniques that use optical sensors benefit from the accessibility provided by the use of smartphones but their nutritional applications thus far are limited to macronutrients. They are also more susceptible to demographic biases in hardware, software, and datasets [110].
5 Discussion and Future Work
This review of the state of the art reveals several limitations and gaps that hinder the assessment of micronutrient status in individuals: (1) the lack of clinical relevance in innovative approaches, (2) the absence of comprehensive assessment techniques, and (3) the deficiency of accessible and noninvasive methods. To address these challenges, we propose high-priority, actionable responses for each limitation (Table 9). These future directions aim to transition micronutrient status assessment from the laboratory to everyday use, enabling valuable micronutritional insights in a manner that is both easily accessible and noninvasive. This approach would empower clinicians to get reliable insights into nutritional status, empower researchers to run cost-effective, larger-scale studies to analyze more data, and empower individuals to manage micronutrient malnutrition effectively. These are necessary steps to not just address micronutrient malnutrition when it occurs, but to be proactive in its prevention.

### v2

@@ 3.6. Biofluid Analysis Limitations
Prediction from clinical health data is able to combine and analyze a large breadth of features, but this data is often insufficient for micronutrition. The lack of micronutrition data availability poses a grand limitation for the ability to make strides in analytic techniques with AI/ML. As a result, Brown et al. (2021) and others urge for more micronutrition data. Additionally, the insights provided by these solutions are limited to a small set of micronutrients and/or indicate only a binary deficiency status, rather than a continuous one.
4. Physiological Sensing for Micronutrient Assessment
Physiological assessment is as important as biofluid analysis, and some argue they should be considered together. Many physiological symptoms of micronutrient deficiency (Tables LABEL:tab:_physio-ws to LABEL:tab:_physio-min) only manifest in severe cases, but the assessment of physiological symptoms remains a critical step in nutritional practice for two central reasons. First, it is useful when estimating the burden of micronutrient malnutrition in a population that is difficult to assess via biofluid methods. Second, physiological assessments provide insights into patient health that are unique and complementary to quantitative, biofluid-based assessments. Though it is valuable, physiological analysis receives significantly less attention in emerging micronutrition research. A comprehensive overview of physical examination in clinical nutrition is out of scope for this review, so we direct the reader to Reber et al. (2019) and Hummell and Cummings (2022) for more information.
Physiological sensing enables automatic detection of symptoms linked to micronutrient deficiency (Section 2.3 and Tables LABEL:tab:_char-ws to LABEL:tab:_char-min), supporting NFPEs. Incorporating metrics like heart rate, activity, sleep, and temperature can enhance personalization and reveal health trends over time. King et al. (2017) outline key criteria for effective wearable monitoring: noninvasive, user-friendly, reliable, and informative. Witt et al. (2019) shows how raw data from sensors like PPG, ECG, accelerometers, EDA, and temperature can offer physiological insights about human physiology relevant to micronutrient imbalance. Yokus and Daniele (2021) also provide valuable design considerations for future wearable micronutrient assessment tools. The remainder of this section highlights innovative methods of optical sensors in assessing physiological signals. While not explicitly wearables-focused, Yokus and Daniele (2021) provide some useful considerations for wearable devices that should be applied to future micronutrient assessing devices.
4.1. Applications of Optical Sensors in Assessing Physiological Signals
@@
McDuff (2021) reviews how camera sensors can be used for noninvasive physiological measurement, which can allow for nutritional insights. The analysis of motion artifacts can reveal minute subtleties in body motion over time that are caused by various physiological mechanisms (e.g. breathing). Also, camera sensors can measure the intensity and wavelength of light absorbed and reflected by our bodies, especially skin (Fig. 4). Differences in measured light over time can be observed and associated with physiological signals (e.g. heartbeat) and status (e.g. low blood oxygen saturation). However, it is important to note that skin melanin content can impact how light is absorbed and reflected (Austin et al., 2021). A failure to properly account for these differences can (and has) resulted in racial biases (for example, pulse oximetry (Sjoding et al., 2020) and wrist-worn heart rate sensors (Koerber et al., 2023)).
@@
Figure 4. As light is applied to the skin, it is absorbed and reflected in different ways depending on its wavelength. Importantly, skin melanin content can impact the absorbance of visible light. For more information, please refer to (Austin et al., 2021). Used with permission from McDuff (2021).
[…]
The figure shows the distribution of different kinds of waves on the electromagnetic spectrum along a horizontal line. On the left are gamma rays, x rays, and ultraviolet light, in the middle of the spectrum is the visible light range (indicated by the classic rainbow), and on the right of the line are radar, microwaves, radio waves, etc. The wavelength of each is also shown along the line. Below this wavelength line, the figure provides a zoomed in image of the visible light spectrum with colors ranging from purple on the left to black on the right. Underneath the spread of these different colors is an image of skin with the stratum corneum at the top (.1 mm thick), the epidermis below it (.1-1 mm thick), followed by the dermis (1-3 mm thick) and including blood capillaries and veins. There are arrows that drop down from each of the colors on the visible light spectrum demonstrating how far they penetrate into the skin. Going from the left we have purple light with the smallest wavelength and the shallowest penetration into the skin. The figure also notes that with less penetration there is less scattering. There is a rather linear increase across the cyan, green, yellow, red and black colored wavelengths penetration and scattering. Black is located on the right and has the greatest penetration and scattering, extending all the way past the dermis.
@@
The type of optical sensor has a large impact on the signals that can be derived from it (McDuff, 2021). RGB cameras are found in most smartphones and, therefore, are the most prevalent. These operate largely in the visible spectrum of light (400 to 700 nm wavelengths), but they can often detect some light in the NIR range. NIR cameras are able to detect light in the 700 to 1000 nm range, and thermal cameras can go fully into the infrared spectrum of 2000 to 14000 nm. Thermal cameras, as their name suggest, can provide unique information over other sensors such as body temperature and sweat gland activation. However, this comes with a higher cost and lower resolution. Finally, multi and hyper-spectral cameras allow for the measurement of multiple wavelengths of light at once. This can also be achieved by combining signals from multiple sensors.
@@
Cameras have been extensively applied to the measurement of physiological vital signs (McDuff, 2021; McDuff and Blackford, 2019; Liu et al., 2023; Nabian et al., 2018; Van Der Kooij and Naber, 2019; Guede-Fernández et al., 2020). One survey describes 5 physiological vital signs where optical sensor measurements been applied: pulmonary activity, EDA, blood oxygen saturation, glucose status, and cardiac activity (McDuff, 2021). PPG signal capture has been achieved with RGB cameras (Van Der Kooij and Naber, 2019; Guede-Fernández et al., 2020). NIR cameras are also useful for PPG in use cases like sleep monitoring that require environments with low visible light (McDuff, 2021). When measuring blood oxygen saturation, multi or hyper-spectral cameras are preferred. Measuring multiple wavelengths of light at once benefits the simultaneous measurement of oxy and deoxy-hemoglobin, and therefore blood oxygen saturation. Recently, Sharma et al. (2023) proposed a smartphone-based hyperspectral imaging platform and used it to identify organic fruits. Optical sensor-based techniques could be use to detect and investigate other physiological symptoms in the fingertip (Hasan et al., 2018; Wang et al., 2016; Umar et al., 2020), mouth (Welikala et al., 2020; Xue et al., 2022), and eyes (Suner et al., 2021; Park et al., 2020; Kumar et al., 2021; Kasiviswanathan et al., 2020).
@@
Optical sensors have been successfully applied to noninvasively assess a major protein, hemoglobin. Hemoglobin allows for oxygen transport in red blood cells and is produced by iron, vitamin B9, and vitamin B12 (Staff, 2023). A deficiency in these micronutrients can result in lowered hemoglobin, which manifests as anemia. A review of the state of the art emphasizes the need for an affordable and accessible method of hemoglobin measurement, which can be realized with commodity smartphone cameras (Hasan et al., 2021). Various works have used smartphone cameras as a way to estimate hemoglobin status noninvasively (Hasan et al., 2021; Wang et al., 2016; Hasan et al., 2018; Suner et al., 2021). They report that PPG signals derived from the fingertip and conjunctiva (skin behind the lower eyelids) under NIR light in 1070 and 850 nm wavelengths contain the most critical features to hemoglobin estimation. Wang et al. (2016) had users place their fingertip directly onto a smartphone camera while it recorded video to determine hemoglobin status. Analysis found that the blue spectra of plasma was the most important for protein composition, and therefore hemoglobin estimation. Classifiers were built to identify hemoglobin status as “normal” or “anemic” in the context of demographic averages. Optimal reference standard blood hemoglobin measurements (Masimo Pronto optical device) were used as ground-truth. Because camera-based assessment can exhibit bias with differences in skin color (see above), the authors carefully considered and reported the demographics of their study participants in addition to controlling for light absorption by skin tissue during experiments. A later paper devised a similar system, applying artificial neural networks (ANNs) and using only the smartphone’s flash as a light source (Hasan et al., 2018). By combining frames within the videos, the researchers were able to identify regions with high variation in predicted levels.
@@
The use of the conjunctiva as an ROI is gaining popularity due to the fact that it, like the fingernail bed and palmar creases, has no melanin and is devoid of “epidermis, dermis or subcutaneous fat which could impede the transmission of light” to deeper vascular layers (Suner et al., 2021). This means the blood vessels are easier to analyze, and there may be less bias due to skin color. These properties have been leveraged by Suner et al. (2021) to estimate hemoglobin concentration and screen for anemia using smartphone images of the conjunctiva. Spectral super-resolution (SSR) has been introduced to measure blood hemoglobin levels (Park et al., 2020). This method is based on a wealth of existing research reconstructing hyperspectral images from RGB signals. Statistical learning was applied to approximate a hyperspectral image of the conjunctiva from a simple smartphone camera image. Using the hyperspectral data, the hemoglobin content in blood can be computed more effectively than the RGB data alone. However, no analysis of skin color differences was conducted.
@@
Techniques that use optical sensors benefit from the accessibility provided by the use of smartphones but their nutritional applications thus far are limited to macronutrients. They are also more susceptible to demographic biases in hardware, software, and datasets (McDuff, 2021).
5. Implications for Future Micronutrient Status Assessment Methods
The state of the art reveals limitations that hinder the assessment of micronutrient status in individuals in the following ways: (1) the lack of clinical relevance in innovative approaches, (2) the absence of comprehensive assessment techniques, and (3) the deficiency of accessible and noninvasive methods. Future work could aim to address these issues while considering real-world integration into clinical and public health settings. Here, the requirements for medical diagnostic tests set by frameworks like REASSURED (Land et al., 2019) are useful as they intersect with and expand beyond accessibility and non-invasiveness. REASSURED stands for Real-time connectivity, Ease of specimen collection, Affordable, Sensitive, Specific, User-friendly (including cultural acceptability), Rapid and robust, Equipment-free/Environmentally friendly, and Deliverable (i.e. accessible) to end-users (Land et al., 2019). First we briefly discuss how emerging methods align with these criteria, summarized in Table 10, then we discuss opportunities to address the aforementioned limitations, summarized in Table 11. We note that definitive conclusions remain challenging due to the early stage of many technologies.
Table 10. Whether Methods Address Proposed and REASSURED (Land et al., 2019) Criteria.
Optimal Reference Standard Assays Electrochemical Spectro-scopic Biofluid Analysis
Criteria - Sweat Colorimetry Multiplexed Sweat Sensors Smartphone-based Commercial Voltammetry Amperometry Impedance - Single Multiple
Clinical relevance X - - X X - - - X X X
Comprehensive assessment - - X - - - - - - - X
Noninvasive - X X - - X X X - X X
Accessible - X X X - - - - - - -
Reproducibility X - - - X - - - X - -
Robustness - - - - - - - - - X X
REASSURED
Real-time connectivity - - X X X - - - - - -
Ease of specimen collection - X X - - - - - - - -
Affordable - X X X - X X X - - -
Sensitive X - - X X - - - X - -
Specific X - - X X - - - X - -
User-friendly - X X - - - - - - X X
Rapid - X X X X X X X X X X
Equipment-free/ Environmentally friendly - X X - - - - - - - -
Deliverable - X X X - - - - - - -
5.1. Evaluating Emerging Methods
Assay-based methods like lateral flow immunoassays offer high sensitivity and specificity due to their targeted biomarker reactions but face issues with batch variability, limited range sensitivity, and environmental instability. Their performance can vary across sample types and demographics. While assay-based methods are familiar to clinicians, they may burden patients due to reliance on blood or urine samples.
Electrochemical methods are promising for PoC use due to speed and simplicity but can be affected by biofluid properties, electrode placement, and demographic differences. They often rely on less validated biosamples, limiting clinical adoption, and may require benchtop equipment, limiting practical deployment in limited resource settings.
Spectroscopy-based methods, such as LC-MS/MS, are the clinical optimal reference standard due to their high reproducibility, but they require complex lab equipment and protocols, limiting field use. Emerging on-body spectroscopic approaches could improve accessibility and cultural acceptability, especially when integrated with smartphones. However, research on noninvasive, accessible spectroscopy for in-body micronutrient assessment remains inadequate.
Biofluid analytic methods, including those applying AI/ML to physiological or multi-modal sensor data, may offer the highest patient acceptability by avoiding additional testing. While potentially fast and accurate, their effectiveness depends heavily on training data quality. Therefore, these approaches face challenges in reproducibility, bias, and scalability, especially across diverse and/or underserved populations. Collecting equitable, high-quality data for training is resource-intensive, and it is likely that effective models will have to be highly specific to the communities they serve, as opposed to being fully generalizable to micronutrient malnutrition globally.
Although emerging methods for accessible and noninvasive micronutrient status assessment have promising capabilities within this set of requirements, few of these approaches are currently standardized and ready for reproducible deployment across diverse geographic settings and clinical populations. The actionable future opportunities we propose in Table 11 aim to transition micronutrient status assessment from the laboratory to everyday use, enabling valuable micronutritional insights in a manner that is both easily accessible and noninvasive.